# The Carbon Footprint Wizard: A Knowledge-Augmented AI Interface for Streamlining Food Carbon Footprint Analysis

### v1

@@ 3.2 Discussion
Second, inconsistencies between the footprints reported in the databases result in considerable impact ranges rather than precise values, as evidenced by the wide min-max intervals (0.13-0.84 kg CO2-eq) presented in the results. This variability is partly due to inherent differences in how each database handles lifecycle phases and system boundaries. Another aspect of data consistency concerns the lack of standardized vocabulary used in the different databases. The semantic matching process, while generally effective, occasionally suggests irrelevant products (such as "tapenade" for olives or "cream sauce with herbs" for oregano), reflecting the inherent challenges of grounding natural language ingredient descriptions into standardized database terminology. While our current methodology employs a human-in-the-loop approach to mitigate this issue (by allowing users to select appropriate matches), it introduces dependency on user expertise and judgment, creating another potential source of variability in the final assessment results.
@@
Finally, the system also demonstrates limitations in handling complex analytical tasks using a single LLM. For instance, the cooking impact estimation occasionally fails to detect implied cooking requirements, as demonstrated by the veggie pizza example, where the system incorrectly indicated "no cooking required" despite pizza preparation typically involving baking. Such commonsense violations can potentially lead to a significant underestimation of the total environmental impact of the recipe, as cooking can contribute substantially to the CF of a meal. Similarly, the generation of follow-up questions sometimes falls short of identifying the most relevant aspects for further exploration, focusing on generic queries rather than insights specific to the analyzed recipe’s unique characteristics. These issues stem from the ambitious scope of responsibilities assigned to a relatively small language model (GPT-4o-mini), including ingredient extraction, impact calculation, cooking estimation, visualization, data preparation, scaling real-world equivalence, and follow-up question generation. The model must simultaneously interpret heterogeneous data formats from multiple databases, perform mathematical operations, generate natural language explanations, and create structured visualization data, all within a single inference pass. This overextension of the model’s capabilities could be addressed in future iterations by agentic AI, which can decompose these tasks into separate, specialized LLM calls that focus on distinct aspects of the analytical pipeline. This approach may involve employing larger and more capable models for the most complex reasoning tasks.
Prompts used to guide the LLMs in the Carbon Footprint Wizard and example outputs corresponding to the case study described in the Results section.
@@
The first author has been supported by a seed grant awarded by the Vrije Universiteit’s Amsterdam Sustainability Institute (ASI) to the second and the third author.

### v2

@@ Discussion
Second, inconsistencies between the footprints reported in the databases result in considerable impact ranges rather than precise values, as evidenced by the wide min-max intervals (0.13-0.84 kg CO2-eq) presented in the results. This variability is partly due to inherent differences in how each database handles lifecycle phases and system boundaries. Another aspect of data consistency concerns the lack of standardized vocabulary used in the different databases. The semantic matching process, while generally effective, occasionally suggests irrelevant products (such as ”tapenade” for olives or ”cream sauce with herbs” for oregano), reflecting the inherent challenges of grounding natural language ingredient descriptions into standardized database terminology. While our current methodology employs a human-in-the-loop approach to mitigate this issue (by allowing users to select appropriate matches), it introduces dependency on user expertise and judgment, creating another potential source of variability in the final assessment results.
@@
Finally, the system also demonstrates limitations in handling complex analytical tasks using a single LLM. For instance, the cooking impact estimation occasionally fails to detect implied cooking requirements, as demonstrated by the veggie pizza example, where the system incorrectly indicated ”no cooking required” despite pizza preparation typically involving baking. Such commonsense violations can potentially lead to a significant underestimation of the total environmental impact of the recipe, as cooking can contribute substantially to the CF of a meal. Similarly, the generation of follow-up questions sometimes falls short of identifying the most relevant aspects for further exploration, focusing on generic queries rather than insights specific to the analyzed recipe’s unique characteristics. These issues stem from the ambitious scope of responsibilities assigned to a relatively small language model (GPT-4o-mini), including ingredient extraction, impact calculation, cooking estimation, visualization, data preparation, scaling real-world equivalence, and follow-up question generation. The model must simultaneously interpret heterogeneous data formats from multiple databases, perform mathematical operations, generate natural language explanations, and create structured visualization data, all within a single inference pass. This overextension of the model’s capabilities could be addressed in future iterations by agentic AI, which can decompose these tasks into separate, specialized LLM calls that focus on distinct aspects of the analytical pipeline. This approach may involve employing larger and more capable models for the most complex reasoning tasks.
Conclusions
This paper introduced an interdisciplinary methodology that bridges state-of-the-art LCA frameworks and databases with the current generation of knowledge-augmented AI. The methodology’s four steps leveraged LLMs for seamless user interaction, fueled by LCA databases with contextual knowledge about food products. The methodology was designed to involve users in the ingredient selection step and to allow users to explore the data further after the summarized report was presented. The methodology is implemented in the publicly available Carbon Footprint Wizard web interface. The initial analysis of the tool’s functionality showcased the power of connecting LCA insights with LLMs, providing an intuitive entry point to complex information for lay users. At the same time, the analysis revealed challenges with missing data, inconsistencies between the databases, and violations of common sense by the LLM. These challenges point to substantial interdisciplinary obstacles that can be fruitfully explored in the future: the engineering and representation of LCA knowledge, the development of LCA data quality frameworks, and ensuring consistency in LLM behavior. Agentic AI frameworks and advanced entity resolution approaches can be employed to address these limitations. Future work should address two limitations of our approach. First, the lack of quantitative validation of our method necessitates benchmarking against expert LCAs, conducting sensitivity/error analyses, and conducting user studies to demonstrate accuracy, robustness, and usability at scale. Second, our three databases differ in geographical scope, system boundaries, and category definitions, leading to wide impact ranges and limited comparability. Thus, future work should revise the selection of LCA databases to account for their differences in coverage and focus.
Acknowledgements
@@
This work has been supported by the seed grant “Knowledge-based support for sustainable diet policies” awarded to the authors by the Vrije Universiteit’s Amsterdam Sustainability Institute (ASI).
Appendix A Prompts Used in the System
Ingredient Extraction Prompt
Result Generation Prompt
Appendix B Example Outputs from Each Step
Below are the example outputs from each step of the environmental impact analysis process for a vegetarian pizza recipe.
User Recipe Input
Ingredient Extraction
Product Matching
A screenshot of the matching step in our interface is shown in Figure 4.
Carbon Footprint Data
A visual summary of the analysis results provided by our interface is shown in Figure 5.
Natural Language Report
An example report in our interface is shown in Figure 3. An example of a follow-up user query and a system response is shown in Figure 6.